If 50g of tetraphosphorus decaoxide is mixed with 50g of water, how much phosphoric acid will be produced?
1
Expert's answer
2018-03-07T13:28:07-0500
P4O10 + 6H2O → 4H3PO4
n( P4O10)=50/284=0.176 mol
n(H2O)=50/18=2.77 mol
n(H2O)>6 n(P4O10)
n (H3PO4)=4 n( P4O10)=4*0.176=0.704 mol
m (H3PO4)=n*M=0.704*98=69g
Answer: 69g